A 2.47micro-F and a 6.38 micro-F capacitor are connected in series across a 40.0-V battery. A 9.80 micro-F capacitor is then connected in parallel across the 2.47micro-F capacitor. Determine the voltage across the 9.80 micro-F capacitor.
Answer #1

let

C1 = 2.47 micro F

C2 = 6.38 micro F

C3 = 9.8 micro F

V_battery = 40 volts


C13 = C1 + C3

= 2.47 + 9.8

= 12.27 micro F

voltage across C3, V3 = V_battery*C2/(C13 + C2)

= 40*6.38/(6.38 + 12.27)

= 13.7 volts <<<<<<<<------------Answer